# Benchmarking XAI Explanations with Human-Aligned Evaluations

### v1

@@ C.3.1 Explanation embeddings
$\phi_{\text{image}}({\bm{e}}_{i}^{j})=\text{CLIP}_{\text{image}}({\bm{x}}_{i}\times{\bm{e}}_{i}^{j})$ (11)
@@
Concerning concept bottleneck explanation, which basically can be interpreted as a dictionary attributing a scalar for each concept. One crucial step is converting CBM activations to CLIP embeddings. We tested two ways do do so:
@@
  • By using a sum of all the CLIP embeddings of text, weightened by its activations:
    $\phi_{text}({\bm{e}}_{i}^{j})=\frac{1}{||{\bm{e}}_{i}^{j}||}\sum_{k}^{N_{k}}{\bm{e}}_{i}^{j}[k]~{}CLIP_{text}(concept_{i}[k])$ (12)
If we use the first solution, there is questions about the number of concepts to keep, that we note as the parameters $N_{top}$. Table 15 presents the influence of $N_{top}$ while 16 presents the influence of diffetents ways to compute the embedings.
@@
Weightened 0.86 $\pm$ 0.06 0.53 $\pm$ 0.05 0.27 $\pm$ 0.28
Sentence 0.85 $\pm$ 0.06 0.55 $\pm$ 0.04 0.28 $\pm$ 0.28 “ ”
Sentence 0.82 $\pm$ 0.08 0.54 $\pm$ 0.06 0.27 $\pm$ 0.28 “The model’s prediction is motivated by the concepts: ”

### v2

@@ D.2 Ablation studies
$\phi_{\text{image}}({\bm{e}}_{i}^{j})=\text{VLM}_{\text{image}}({\bm{x}}_{i}\times{\bm{e}}_{i}^{j})$ (14)
@@
Concerning concept bottleneck explanation, which basically can be interpreted as a dictionary attributing a scalar for each concept. One crucial step is converting CBM activations to VLM embeddings. We tested two ways do do so:
@@
  • By using a sum of all the VLM embeddings of text, weightened by its activations:
[…]
If we use the first solution, there are questions about the number of concepts to keep, that we note as the parameters $N_{top}$. Table 19 presents the influence of $N_{top}$ while 20 presents the influence of differents ways to compute the embeddings. Analysis of the results indicates that an adequate number of concepts enhances ranking performance, which eventually plateaus, thereby informing the selection of $N_{top}=15$. Regarding the embedding procedure, optimal outcomes were achieved by encoding the top concepts as a sentence. The inclusion of a template demonstrated negligible influence on overall performance metrics.
@@
Weighted 1.212 $\pm$ 0.288 0.392 $\pm$ 0.077 0.425 $\pm$ 0.082
Sentence 0.989 $\pm$ 0.113 0.471 $\pm$ 0.055 0.501 $\pm$ 0.052
Sentence 0.986 $\pm$ 0.119 0.469 $\pm$ 0.061 0.504 $\pm$ 0.055
“Concepts in explanation:”
Appendix E E. PASTA-score: Additional Examples
Table 21: PASTA scores corresponding to different explanations. Dataset: catsdogscars. Label: dog
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
P (Q1) 2.01 2.04 3.94 3.01 3.81 3.97 2.14
P (Q2) 1.90 1.82 3.31 2.80 3.57 3.37 1.70
P (Q3) 2.93 2.58 3.65 4.06 4.02 3.65 2.83
P (Q4) 3.04 2.95 3.84 4.09 4.14 3.84 2.75
Table 22: PASTA scores corresponding to different explanations. Dataset: MonumAI. Label: Baroque
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
P (Q1) 2.45 1.01 3.06 4.00 1.00 1.67 3.22
P (Q2) 2.51 1.26 2.99 3.48 1.57 1.52 3.51
P (Q3) 3.67 1.52 3.51 3.71 2.05 1.54 3.69
P (Q4) 3.35 1.33 3.45 3.89 1.33 1.59 3.79
Table 23: PASTA scores corresponding to different explanations. Dataset: COCO. Label: transportation
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
P (Q1) 1.21 3.24 2.46 3.44 2.01 2.80 2.40
P (Q2) 1.61 3.22 2.45 3.36 2.23 2.75 2.37
P (Q3) 1.43 3.94 3.45 3.78 3.01 3.69 3.41
P (Q4) 1.59 3.65 3.20 3.50 2.96 3.25 3.28
Table 24: PASTA scores corresponding to different explanations. Dataset: PascalPart. Label: cat
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
P (Q1) 2.02 1.68 3.90 2.51 3.93 2.50 3.30
P (Q2) 2.53 1.71 3.91 2.55 3.90 2.60 3.42
P (Q3) 2.88 2.77 3.78 3.22 3.94 3.30 3.80
P (Q4) 3.06 2.82 3.78 3.27 3.82 3.26 3.82
Table 25: PASTA scores corresponding to different explanations. Dataset: PascalPart. Label: person
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
P (Q1) 3.68 3.55 3.69 3.12 1.35 3.71 1.93
P (Q2) 3.46 3.19 3.16 3.07 1.74 3.16 1.81
P (Q3) 3.50 3.64 3.64 3.02 1.77 3.59 2.56
P (Q4) 4.23 4.09 4.05 3.82 1.29 4.04 1.91
Table 26: PASTA scores corresponding to different explanations. Dataset: COCO. Label: cultural
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
P (Q1) 2.65 2.27 2.78 1.21 2.96 2.75 1.45
P (Q2) 2.21 1.73 2.31 1.64 2.82 2.14 1.83
P (Q3) 2.40 2.17 3.05 1.37 3.25 2.49 1.72
P (Q4) 2.49 2.20 2.84 1.30 3.06 2.75 1.92
Table 27: PASTA scores corresponding to different explanations. Dataset: catsdogscars. Label: car
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
P (Q1) 3.20 2.67 4.80 2.10 2.74 4.80 1.50
P (Q2) 3.18 2.60 4.69 1.53 2.37 4.69 1.64
P (Q3) 2.98 2.54 4.30 2.55 3.17 4.30 2.16
P (Q4) 3.34 2.47 4.06 2.51 3.01 4.06 1.72
Table 28: PASTA scores corresponding to different explanations. Dataset: PascalPart. Label: motorbike
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
P (Q1) 1.26 1.42 3.38 2.43 1.10 1.00 3.26
P (Q2) 1.59 1.61 3.35 2.22 1.46 1.52 3.07
P (Q3) 2.18 1.43 4.41 3.31 1.96 1.77 4.18
P (Q4) 2.15 2.10 4.22 3.47 1.45 2.24 4.08
Table 29: PASTA scores corresponding to different explanations. Dataset: MonumAI. Label: Gothic
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
P (Q1) 2.74 3.75 3.38 3.00 3.97 2.88 3.69
P (Q2) 2.51 3.43 3.13 2.86 3.91 2.79 3.44
P (Q3) 3.21 3.90 3.58 3.43 4.37 3.70 3.98
P (Q4) 3.18 3.85 3.51 3.38 4.25 3.52 3.84